An infinitely long wire is kept along z=-infinity axis from z=+infinity, having uniform linear charge density 10/9nC/m. The electric field at a point ( 6cm, 8cm, 10cm ) will be 

Grade:12

2 Answers

Khimraj
4 years ago
E= (lambda)/(2pi epsilon_(0)d)= (2K lambda)/(d)` 
`=(2xx9xx10^(9)xx(10//9)xx10^(-9))/(10xx10^(-2))=200 V//m` 
As field is radial , out, there is no component of it in `Z` direction.  
`vec(E )=(200 cos 53^(@) hat (i)+200 sin 53^(@)hat(j))V//m` 
`=(200xx(3)/(5)hat(i)+200xx(4)/(5)hat(j))= (120hat(j)+160hat(j))V//m`
Khimraj
4 years ago
electric field due to infinite rod is equal to (charge density)/(2πεr) where r is dist of point from rod.....here sincerod is along z axis therefore r will be 10 cm.....
therefore (10/9)*10^-9/2πε10*10^-2
=200k V/m